OpenStudy (anonymous):

evaluate using product quotient chain sec(2theta)tan^2(theta)

OpenStudy (anonymous):

This is tough. So this combines both product and chain rules. If you have a product of functions f*g, then the derivative of that product is f’ * g + g’ * f. So if f = sec(2x) and g = tan^2 (x), then f’ = [2] [sec (2x) * tan (2x)] (where the 2 comes from the chain rule) and g’ = 2 * tan x [again chain rule]. So f’ * g + g’ * f = ([2] [sec (2x) * tan (2x)])* (tan^2 (x)) + [2 * tan x] * [sec(2x)].